Problem 3 (10 points): FIN300 Inc. recently issued non-callable, semi-annually paying bonds that mature in 15 years. They have a par value of $1,000 and an annual coupon of 5.5%. If the current market interest rate is 6.0%, at what price should the bonds sell?

Problem 4 (15 points): FIN300 Ltd.s outstanding bonds have a S1,000 par value, and they mature in 20 years. Their nominal yield to maturity is 9.00%, they pay interest semiannually, and they sell at a price of $985. What is the bonds nominal coupon interest rate?

Can you also include the formulas as well
Answer #1
Q3.
Annuity PVF at 3% for 30 periods 19.60044
PVF at 3% for 30th period 0.411987
Par value of bonds 1000
Semi Annual Cash interest(1000*5.50%*6/12) 27.5
Present value of Maturity 411.987
Present value f Interest 539.0121
Price of Bonds 951
Q4.
Annuity PVF at 4.58% for 40 periods 18.40158
PVF at 4.58% for 40th period 0.171929
Par value of bonds 1000
Semi annual cash interest    x
Present value of Maturity 171.929
Present value f Interest 18.40158 x
Price of Bonds 171.929+18.40158x
As per given in Question, the price of bonds is $ 985. Hence, the equation is:
985 = 171.929+ 18.40158 x
x= 44.185
Semi annual interest = 44.185
Annualised Interest = 44.185*2 = 88.37
Annual coupon interest rate = 88.37 /1000*100 = 8.837%